OpenStudy (anonymous):

Solve the system. y = x + 2 and y = -x + 0 (1, -1) (-1, 1) (1, 3) (-1, 3) no solution

OpenStudy (lgbasallote):

y = x+2 y = -x +0 add them,... 2y = 2 y = 1 x = -1